An engineering technology group just purchased new CAD software for $5000 now and annual payments of $500 per year for 8 years starting 4 years from now for annual upgrades. What is the present worth of the payments if the interest rate is 8% per year?

1
Expert's answer
2022-01-19T10:19:07-0500

"P'_A" is located in year 3

"P_T= P_0+P_A"

=5000+500(P/A,8%,8)(P/F, 8%,4)

=5000+500(5.7466)(0.730)=

=5000+2097.509= 7,097.509